OpenStudy (anonymous):

Solve the following system of equations for x. - 3x + 7y = 2 - 2x - 3y = - 4

OpenStudy (anonymous):

@whpalmer4

OpenStudy (neer2890):

multiply equation 1 by 2 and equation 2 by3 you got -6x+14y=4 -6x-9y=-12 now subtract 1 from 2 23y=16 y=16/23 put this value of y in any equation and find x

OpenStudy (anonymous):

Oh lemme do that

OpenStudy (anonymous):

-3x+4 20/23 =2

OpenStudy (neer2890):

-3x+7*16/23=2

OpenStudy (anonymous):

x=22/23